Systems Staff Meeting Notes 12/20/99

Systems Staff Meeting Notes
December 20, 1999

Present: Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Hilton Soto turns 25 today!
Ongoing Systems Updates

A. Orders
Y2K Upgrade for Novell server

a) Software upgrade - theoretically ready to go
b) Hardware has arrived, but is missing one 128 Simm. It should still work with existing memory. The missing Simm was supposed to be overnighted last Wednesday. Will investigate.
Equipment Received/Installed/Repaired
  1. Security - general panic regarding games viruses - Frogapault & Elf Bowling. The threat is fraudulent; Bill has posted as much to the Liaisons.
  2. There has been increasing incidents of SPAM to Librarians - lots of solicitations from booksellers. As a rule, discard the message. If it recurs, don't respond to the "take me off the list" order (can make spamming worse) - just let Systems know. Sometimes a third party removal is more effective.
Networking
  1. Change in Core Internet POP configuration - public sites and academic offices have changed to Internet addresses. The NAT has been pulled off current setup and put on to separate hardware. We're now showing a drop in activity and expect to have less afternoon congestion. Watch for calls reporting an inability to reach the Library web pages - they may be on private IP, indicating the new translator could be failing.
  2. WebLuis - NERSP machine at FCLA had problems over the weekend.
  3. Smathers network reconfiguration - may occur in January.
  4. NT5 replacement - switchover should occur this afternoon in MSL. This should give them more muscle.
  5. Resource Services - SNA timeout problem - 3270 sessions used for the tech side of Notis have been habitually going down, some older versions have an interactive effect with CLARR. We would like to isolate the problem by changing the offending work stations to the network switch from the HUB. It may or may not rectify the problem. If successful, we'll need to accelerate the switch installations in that area.
  6. West wiring - we received quotes from Digital Design, but following a meeting with John Humphrey, Will learned that they are going to be rewiring West for telephone installation. We will piggyback our network wiring on to that project. Two Com Closets will be installed on each floor, fiber will be run parallel to the stacks, horizontal between the floors. This will require much planning. Systems found out about the phone rewiring project accidentally - we would have wasted lots of money if the current wiring estimates had been sent through.
  7. Preservation, Norman Hall wiring - quotes have been returned, need to have them signed by Martha. Wiring in Norman is for the SNA terminal replacements, in Preservation it will be for the new room, 203A.
  8. Removal of SNA terminals - we'll start deactivating them next week, taking the modems offline.
Software Projects Status
  1. Wiring program - Ying has finished the application file on the server, need to install on Bobbie's machine for testing.
  2. Forms to mail - waiting on server.
  3. OCLC tape conversions - Waiting on NERDC - Winston talked to Jim Arnette (he's been busy) and Sharon Wheeler returns today.
  4. Trouble Ticket Viewer - awaiting feedback from select liaisons, working on setting permissions.
  5. Ed Teague has requested a searchable database of undergrad Architecture students' projects. Cataloging Department will not include these in the library catalog. Hilton created the database in one day. Bill notes that we need to resolve what does & does not go into the catalog. Ed's argument was that this is not a thesis or dissertation and would not be catalogued. Ed will be allowed to edit the database, however, maintenance could become a problem. It really should be in NOTIS, Bill will discuss this topic with Martha.
  6. Hilton is looking at the first module for the Q&A replacement database for Preservation.
DP Coordinator
  1. New student has been hired for Spring term, Melissa Lenhard.
  2. Peter McKay has given Debra the word that we can return the stockpile of Compustat tapes since patrons will now access that information via CRSP Link on the Business school network. Peter will notify Compustat to change the order so that we no longer receive tape shipments.
Calendar
  1. Rich out December 20 - 29.
  2. Bobbie out December 22-24.
  3. Winston out December 27.
  4. Will out December 28-29.
  5. New Year Holiday - Friday, Dec. 31.
  6. Everett out Jan. 6-12.
  7. Debra out Jan. 20-27.
  8. Everett out Jan. 28-31.


Back to the Systems Homepage

Last updated December 27, 1999
by Debra Harris
debharr@mail.uflib.ufl.edu